# VisPath: Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization

### v1

@@ Abstract
Unprecedented breakthroughs in Large Language Models (LLMs) has amplified its penetration into application of automated visualization code generation. Few-shot prompting and query expansion techniques have notably enhanced data visualization performance, however, still fail to overcome ambiguity and complexity of natural language queries - imposing an inherent burden for manual human intervention. To mitigate such limitations, we propose a holistic framework VisPath : A Multi-Path Reasoning and Feedback-Driven Optimization Framework for Visualization Code Generation, which systematically enhances code quality through structured reasoning and refinement. VisPath is a multi-stage framework, specially designed to handle underspecified queries. To generate a robust final visualization code, it first utilizes initial query to generate diverse reformulated queries via Chain-of-Thought (CoT) prompting, each representing a distinct reasoning path. Refined queries are used to produce candidate visualization scripts, consequently executed to generate multiple images. Comprehensively assessing correctness and quality of outputs, VisPath generates feedback for each image, which are then fed to aggregation module to generate optimal result. Extensive experiments on benchmarks including MatPlotBench and the Qwen-Agent Code Interpreter Benchmark show that VisPath significantly outperforms state-of-the-art (SOTA) methods, increased up to average 17%, offering a more reliable solution for AI-driven visualization code generation.
VisPath: Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization
Wonduk Seo1thanks: Equal contribution.   Seungyong Lee2*   Daye Kang1,3   Zonghao Yuan4   Seunghyun Lee1thanks: Corresponding author. 1Enhans.ai   2Unaffiliated   3KAIST   4Tsinghua University 1{seowonduk,seunghyun,daye}@enhans.ai 2leesy7197@khu.ac.kr   4yzh23@tsinghua.edu.cn
@@ 1 Introduction
Data visualization has long been an essential tool in data analysis and scientific research, enabling users to uncover patterns and relationships in complex datasets Vondrick et al. (2013); Demiralp et al. (2017); Unwin (2020); Li et al. (2024a). Traditionally, creating visualizations requires manually writing code using libraries such as Matplotlib, Seaborn, or D3.js Barrett et al. (2005); Bisong and Bisong (2019); Zhu (2013). This approach demands programming expertise and significant effort to craft effective visual representations, which can be a barrier for many users Bresciani and Eppler (2015); Saket et al. (2018); Sharif et al. (2024). As datasets continue to grow in size and complexity, researchers have explored ways to automate visualization generation, aiming to make the process more efficient and accessible Wang et al. (2015); Dibia and Demiralp (2019); Qian et al. (2021).
@@
More recently, LLM-based visualization frameworks such as Chat2VIS Maddigan and Susnjak (2023) and MatPlotAgent Yang et al. (2024) have been introduced to improve automated visualization code generation. Specifically, Chat2VIS follows a prefix-based approach, guiding LLMs to generate visualization code consistently; and MatPlotAgent expands the query before code generation. However, these methods face several limitations: ① both generate code in a single-path manner, limiting exploration of alternative solutions and unable to fix-out when caught in misleading bugs; ② both rely on predefined structures or examples which restrict adaptability to ambiguous or unconventional user queries. ③ both approaches encapsulate limitation in their inability to aggregate and synthesize multi-dimensional feedback. Without a mechanism to retrieve outputs that reflect diverse possibilities, they struggle to capture intricate details, ultimately limiting the precision and adaptability of the generated visualizations.
@@
To address these limitations, we introduce VisPath: A Branch Exploration Framework for Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization, a transformative approach that redefines how visualization code is generated. Traditional single-pass methods may seem efficient, but they often fall short of delivering the depth and precision users truly need. They generate code quickly, but struggle to capture the intricate details that make a visualization not just functional but meaningful. VisPath challenges this limitation by incorporating Multi-Path Reasoning and Feedback-Driven Optimization, systematically exploring multiple interpretative pathways to construct a more accurate, context-aware, and fully executable visualization.
@@
Rather than simply translating user input into code, VisPath ensures that every critical aspect—both explicitly stated and implicitly necessary—is carefully considered, creating a visualization that is not only correct but insightful. At its core, it generates multiple reasoning paths that analyze the user’s intent from different perspectives, producing structured blueprints that are then transformed into visualization scripts through Chain-of-Thought (CoT) prompting. These multiple candidates are evaluated using a Vision-Language Model (VLM) to assess accuracy, clarity, and alignment with the intended message. The results are then refined through an Aggregation Module, optimizing the final output for both reliability and impact. By shifting from a single-pass, one-size-fits-all approach to a dynamic, multi-layered process, VisPath sets a new standard for visualization generation—one that is not just about creating code, but about ensuring data is represented in its most clear, meaningful, and compelling form.
@@
Extensive experiments on benchmark datasets, including MatPlotBench Yang et al. (2024) and Qwen-Agent Code Interpreter Benchmark [^1] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/codeinterpreter/README.md, demonstrate VisPath’s superiority over the state-of-the-art (SOTA) visualization generation methods. By systematically generating and evaluating multiple reasoning paths and leveraging iterative feedback aggregation, VisPath significantly enhances the accuracy, robustness against underspecified queries, and adaptability to diverse user intents. Our investigations demonstrate its ability to capture nuanced user intents, improve execution reliability, and minimize errors, making visualization code generation more accessible and effective for domains such as business intelligence, scientific research, and automated reporting.
@@ 2 Related Work
Building on these machine learning strategies, numerous studies have utilized Large Language Models (LLMs) to further enhance system flexibility. Recent frameworks such as Chat2VIS Maddigan and Susnjak (2023) and Prompt4Vis Li et al. (2024b) utilize few-shot learning or query expansion to refine user queries, subsequently generating Python visualization scripts through instruction-based prompting. More recent approaches, such as MatPlotAgent Yang et al. (2024) and PlotGen Goswami et al. (2025), extend these frameworks by integrating a vision-language feedback model to iteratively optimize the final code based on evaluations of the rendered visualizations. The aforementioned approaches often struggle to effectively capture user intent in complex visualization tasks. By committing to a single reasoning trajectory, they may produce code that is syntactically correct yet semantically misaligned with user expectations, requiring extensive manual adjustments. This challenge is particularly pronounced when user input is ambiguous or underspecified, leading to an iterative cycle of prompt refinement and code modification—ironically undermining the intended efficiency of automation. To address these limitations, we introduce VisPath, a novel framework that integrates multi-path reasoning with feedback from Vision-Language Models (VLMs) to enhance visualization code generation. Compared to conventional methods constrained by a single reasoning path, VisPath dynamically explores multiple solution pathways, improving interpretability, minimizing manual intervention, and adapting more effectively to diverse visualization requirements.
@@ 3 Methodology
We introduce VisPath, a framework for robust visualization code generation that leverages diverse reasoning and visual feedback. VisPath is built on three core components: (1) Multi-Path Query Expansion, which generates multiple reasoning paths informed by the dataset description; (2) Code Generation from Expanded Queries, which synthesizes candidate visualization scripts via Chain-of-Thought (CoT) prompting while grounding them in the actual data context; and (3) Feedback-Driven Code Optimization, where a Vision-Language Model (VLM) evaluates and refines the outputs to ensure generation robust visualization code. An overview of this process is shown in Figure 2.
@@ 3.1 Multi-Path Generation
One of the biggest pitfalls in visualization code generation is rigid interpretation, a single query can have multiple valid visual representations depending on its dataset structure. VisPath mitigates this by generating multiple extended queries within a single interaction. Given a user query $Q$ and a corresponding dataset description $D$, a Multi-Path Agent is employed to expand the query into $K$ distinct reasoning pathways:
@@
where $f_{\_}{\text{mpa}}$ denotes the function of the Multi-Path Agent implemented via an LLM. The dataset description $D$ plays a crucial core in shaping these reasoning paths by providing contextual information about variable types, inherent relationships, and the suitability of different chart types for a more grounded interpretation of the query. Each $R_{\_}i$ serves as a detailed logical blueprint outlining one possible approach to fulfill the visualization request. This design ensures that our framework effectively considers a broad range of potential interpretations, thereby increasing the quality of reasoning and the likelihood of capturing the true user intent—even when queries are ambiguous or underspecified.
@@ 3.2 Code Generation from Reasoning Paths
Once diverse reasoning paths are established, the next stage involves translating each path into executable Python scripts. For every reasoning path $R_{\_}i$ generated in Equation (1), a dedicated Code Generation LLM produces the corresponding visualization code using chain-of-thought (CoT) prompting:
@@
where $f_{\_}{\text{code}}$ represents the code generation function. Unlike naive code generation approaches, here, the dataset description $D$ is explicitly provided to ground the generated code in the actual data context, ensuring that variable names, data types, and visualization parameters align correctly with the underlying data attributes. The generated code $C_{\_}i$ is then executed to render a candidate visualization:
@@
with $f_{\_}{\text{exec}}$ serving as the code execution function. By executing the code directly, we ensure that the visualization accurately reflects the intended operations without reintroducing the dataset description $D$ at this stage. In practice, some generated codes may not be executable. Rather than engaging in an explicit debugging loop, we record the execution status as a binary executability indicator:
@@
The result $Z_{\_}i$ (either the rendered visualization or the error message) is then provided, along with $C_{\_}i$ and the original query $Q$, to the feedback model in the subsequent stage.
@@ 3.3 Feedback-Driven Code Optimization
Most code generation frameworks focus merely on syntactically correct scripts, but our framework goes further. As final stage, VisPath synthesizes the most robust and accurate visualization code by leveraging both the executability information and structured visual feedback. A Vision-Language Model (VLM) is employed to analyze each candidate by evaluating the initial query $Q$, the generated code $C_{\_}i$, and the routed output $Z_{\_}i$. This evaluation is formalized as:
@@
Leveraging the collective code-feedback pairs along with the original query $Q$ and the dataset description $D$, an Integration Module synthesizes the final, refined visualization code:
@@
where $C^{*}$ represents the optimized visualization code and $f_{\_}{\text{integrate}}$ denotes the function that aggregates the strengths of each candidate code alongside its corresponding feedback. This formulation ensures that the final code is not only constructed based on the insights extracted from the candidate outputs but is also meticulously aligned with the original user query and the provided dataset description. The process harnesses the reasoning capabilities of LLM to systematically evaluate the strengths and pinpoint potential weaknesses across all candidate solutions, including those that may initially be non-executable. By synthesizing the most optimal elements from each candidate, the framework generates a final code that best captures the user’s intent while maintaining high standards of reliability, robustness, and execution quality. A step-by-step breakdown of this process is detailed in Algorithm 1 below:
@@
2:Final visualization code $C^{*}$
3:// Step 1: Multi-Path Query Expansion
4:$\{R_{\_}1,R_{\_}2,\dots,R_{\_}K\}\leftarrow f_{\_}{\text{mpa}}(Q,D)$triangleright Generate $K$ distinct reasoning paths; each $R_{\_}i$ outlines a potential interpretation of $Q$ given $D$.
5:// Step 2: Code Generation from Reasoning Paths
6:for $i=1$ to $K$ do
7:     $C_{\_}i\leftarrow f_{\_}{\text{code}}(D,R_{\_}i)$triangleright Generate candidate visualization code $C_{\_}i$ based on reasoning path $R_{\_}i$.
8:     $V_{\_}i\leftarrow f_{\_}{\text{exec}}(C_{\_}i)$triangleright Execute $C_{\_}i$ to obtain candidate visualization $V_{\_}i$.
9:     $\epsilon_{\_}i\leftarrow\begin{cases}1,&\text{if }C_{\_}i\text{ executes successfully}\\ 0,&\text{otherwise}\end{cases}$triangleright Record executability indicator $\epsilon_{\_}i$ for $C_{\_}i$.
10:     $Z_{\_}i\leftarrow\begin{cases}\text{plot image}(V_{\_}i),&\text{if }\epsilon_{\_}i=1\\ \text{error message from }C_{\_}i,&\text{if }\epsilon_{\_}i=0\end{cases}$triangleright Route output: either the rendered visualization or the error message.
11:end for
12:// Step 3: Feedback-Driven Code Optimization
13:for $i=1$ to $K$ do
14:     $F_{\_}i\leftarrow f_{\_}{\text{feedback}}(Q,C_{\_}i,Z_{\_}i)$triangleright Obtain structured feedback $F_{\_}i$ on candidate code $C_{\_}i$ using the routed output $Z_{\_}i$.
15:     $S_{\_}i\leftarrow(C_{\_}i,F_{\_}i)$triangleright Form a unified tuple $S_{\_}i$ combining $C_{\_}i$, $F_{\_}i$.
16:end for
17:$C^{*}\leftarrow f_{\_}{\text{integrate}}\bigl{(}Q,D,\{S_{\_}1,S_{\_}2,\dots,S_{\_}K\}\bigr{)}$triangleright Aggregate the code-feedback pairs along with $Q$ and $D$ to synthesize the final visualization code $C^{*}$.
18:return $C^{*}$

### v2

# Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization

@@ Abstract
Rapid advancements in Large Language Models (LLMs) have accelerated their integration into automated visualization code generation applications. Despite advancements through few-shot prompting and query expansion, existing methods remain limited in handling ambiguous and complex queries, thereby requiring manual intervention. To overcome these limitations, we propose VisPath: a Multi-Path Reasoning and Feedback-Driven Optimization Framework for Visualization Code Generation. VisPath handles underspecified queries through structured, multi-stage processing. It begins by reformulating the user input via Chain-of-Thought (CoT) prompting, which refers to the initial query while generating multiple extended queries in parallel, enabling the LLM to capture diverse interpretations of the user intent. These queries then generate candidate visualization scripts, which are executed to produce diverse images. By assessing the visual quality and correctness of each output, VisPath generates targeted feedback that is aggregated to synthesize an optimal final result. Extensive experiments on widely-used benchmarks including MatPlotBench and the Qwen-Agent Code Interpreter Benchmark show that VisPath outperforms state-of-the-art methods, offering a more reliable solution for AI-driven visualization code generation.
Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization
Wonduk Seo1,2   Seungyong Lee   Daye Kang1,3 Hyunjin An1   Zonghao Yuan4   Seunghyun Lee1thanks: Corresponding author. 1Enhans   2Peking University   3KAIST   4Tsinghua University 1{wonduk,daye,hyunjin,seunghyun}@enhans.ai
@@ 1 Introduction
Data visualization has long been an essential tool in data analysis and scientific research, enabling users to uncover patterns and relationships in complex datasets Vondrick et al. (2013); Demiralp et al. (2017); Unwin (2020); Li et al. (2024a). Traditionally, creating visualizations requires manually writing code using libraries such as Matplotlib, Seaborn, or D3.js Barrett et al. (2005); Bisong and Bisong (2019); Zhu (2013). These approaches demand programming expertise and significant effort to craft effective visual representations, which can be a barrier for many users Bresciani and Eppler (2015); Saket et al. (2018); Sharif et al. (2024). As datasets continue to grow in size and complexity, researchers have explored ways to automate visualization generation, aiming to make the process more efficient and accessible Wang et al. (2015); Dibia and Demiralp (2019); Qian et al. (2021).
@@
More recently, LLM-based visualization frameworks such as Chat2VIS Maddigan and Susnjak (2023) and MatPlotAgent Yang et al. (2024) have been introduced to improve automated visualization code generation. Specifically, Chat2VIS follows a prefix-based approach, guiding LLMs to generate visualization code consistently; and MatPlotAgent expands the query before code generation. However, these methods face several limitations: (1) they generate code in a single-path manner, limiting exploration of alternative solutions and unable to fix-out when caught in misleading bugs; (2) they rely on predefined structures or examples which restrict adaptability to ambiguous or unconventional user queries; (3) they encapsulate limitation in their inability to aggregate and synthesize multi-dimensional feedback. Without a mechanism to retrieve outputs that reflect diverse possibilities, they face difficulties in capturing the intricate details required for visualizations that are both functionally precise and contextually relevant.
@@
To address these limitations, we introduce VisPath: A Branch Exploration Framework for Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization, a novel approach that redefines how visualization code is generated, as illustrated in Figure 1. Traditional methods often fall short of delivering the depth and precision users truly need, struggle to capture the intricate details that make a visualization not just functional but meaningful. VisPath challenges this limitation by incorporating Multi-Path Reasoning and Feedback-Driven Optimization, systematically exploring multiple interpretative pathways to construct a more accurate, context-aware, and fully executable visualization.
@@
Rather than simply translating user input into code, VisPath ensures that every critical aspect, both explicitly stated and implicitly necessary, is carefully considered, creating a visualization that is not only correct but insightful. At its core, it generates multiple reasoning paths that analyze the user’s intent from different perspectives, producing structured blueprints that are then transformed into visualization scripts through Chain-of-Thought (CoT) prompting. These multiple candidates are evaluated using a Vision-Language Model (VLM) to assess accuracy, clarity, and alignment with the intended message. The results are then refined through an Aggregation Module, optimizing the final output for both reliability and impact.
@@
Extensive experiments on benchmark datasets, including MatPlotBench Yang et al. (2024) and Qwen-Agent Code Interpreter Benchmark [^1] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/codeinterpreter/README.md, demonstrate that VisPath notably outperforms state-of-the-art methods across all evaluation metrics. In addition, ablation studies confirm that VisPath’s performance gains primarily stem from its Multi-Path Reasoning mechanism and feedback-driven optimization. By systematically generating and evaluating multiple reasoning paths and leveraging feedback aggregation, VisPath significantly enhances the accuracy, robustness against underspecified queries, and adaptability to diverse user intents. Our investigations demonstrate its ability to capture nuanced user intents, improve execution reliability, and minimize errors, making visualization code generation more accessible and effective for domains such as business intelligence, scientific research, and automated reporting.
@@ 2 Related Work
Building on these machine learning strategies, numerous studies have utilized LLMs to further enhance system flexibility. Recent frameworks such as Chat2VIS Maddigan and Susnjak (2023) and Prompt4Vis Li et al. (2024b) utilize few-shot learning or query expansion to refine user queries, subsequently generating Python visualization scripts through instruction-based prompting. More recent approaches, such as MatPlotAgent Yang et al. (2024) and PlotGen Goswami et al. (2025), extend these frameworks by integrating a vision-language feedback model to iteratively optimize the final code based on evaluations of the rendered visualizations. The aforementioned approaches often struggle to effectively capture user intent in complex visualization tasks. By committing to a single reasoning trajectory, they may produce code that is syntactically correct yet semantically misaligned with user expectations, requiring extensive manual adjustments. This challenge is particularly pronounced when user input is ambiguous or underspecified, leading to an iterative cycle of prompt refinement and code modification, consequently limiting the intended efficiency of automation. To address these limitations, we introduce VisPath, a novel framework that integrates Multi-Path Reasoning with feedback from VLMs to enhance visualization code generation.
@@ 3 Methodology
We introduce VisPath, a framework for robust visualization code generation that leverages diverse reasoning and visual feedback. VisPath is built on three core components: (1) Multi-Path Query Expansion, which generates multiple reasoning paths informed by the dataset description; (2) Code Generation from Expanded Queries, which synthesizes candidate visualization scripts via Chain-of-Thought (CoT) prompting while grounding them in the actual data context; and (3) Feedback-Driven Code Optimization, where a VLM evaluates and refines the outputs to ensure generation robust visualization code. An overview of this process is shown in Figure 2.
@@ 3.1 Multi-Path Generation
The potential for rigid interpretation is a major limitation in visualization code generation, a single query can have multiple valid representations depending on its dataset structure. VisPath mitigates this by generating multiple extended queries within a single interaction. Given a user query $Q$ and a corresponding dataset description $D$, a Multi-Path Agent is employed to expand the query into $K$ distinct reasoning pathways:
@@
where $f_{\_}{\text{mpa}}$ denotes the function of the Multi-Path Agent implemented via an LLM. The dataset description $D$ plays a crucial core in shaping these reasoning paths by providing contextual information about variable types, inherent relationships, and the suitability of different chart types for a more grounded interpretation of the query.
Each $R_{\_}i$ serves as a detailed logical blueprint outlining one possible approach to fulfill the visualization request. The purpose of generating multiple reasoning paths is not simply to increase their quantity, but to ensure interpretive diversity. Vispath mitigates the risk of depending on a single, potentially incorrect assumption by intentionally exploring alternative and context-aware reasoning strategies. This design ensures that our framework effectively considers a broad range of potential interpretations, thereby increasing the quality of reasoning and the likelihood of capturing the true user intent even when queries are ambiguous or underspecified.
@@ 3.2 Code Generation from Reasoning Paths
Once diverse reasoning paths are established, the subsequent stage involves translating each path into executable Python scripts. For each reasoning path $R_{\_}i$ generated in Equation 1, a dedicated Code Generation LLM produces the corresponding visualization code using Chain-of-Thought (CoT) prompting:
@@
where $f_{\_}{\text{code}}$ represents the code generation function. The dataset description $D$ is explicitly provided to ground the generated code in the actual data context, ensuring that variable names, data types, and visualization parameters align correctly with the underlying data attributes. The generated code $C_{\_}i$ is then executed to render a candidate visualization:
@@
with $f_{\_}{\text{exec}}$ serving as the code execution function. By executing the code directly, we ensure that the visualization accurately reflects the intended operations without reintroducing the dataset description $D$ at this stage, as the code has already been conditioned on $D$ during its generation.
In practice, some generated codes may not be executable. Rather than engaging in an explicit debugging loop, we record the execution status as a binary executability indicator:
@@
The result $Z_{\_}i$ (either the rendered visualization or the error message) is then provided, along with $C_{\_}i$ and the initial query $Q$, to the feedback model in the subsequent stage.
@@ 3.3 Feedback-Driven Code Optimization
While most code generation frameworks primarily focus on producing syntactically correct scripts, our framework extends beyond this by incorporating an additional mechanism. As final stage, VisPath synthesizes the most robust and accurate visualization code by leveraging both the executability information and structured visual feedback. A VLM is employed to analyze each candidate by evaluating the initial query $Q$, the generated code $C_{\_}i$, and the routed output $Z_{\_}i$. This evaluation is formalized as:
@@
Leveraging the collective code-feedback pairs along with the initial query $Q$ and the dataset description $D$, an Integration Module synthesizes the final, refined visualization code:
@@
where $C^{*}$ represents the optimized visualization code and $f_{\_}{\text{integrate}}$ denotes the function that aggregates the strengths of each candidate code alongside its corresponding feedback. This formulation ensures that the final code is not only constructed based on the insights extracted from the candidate outputs but is also meticulously aligned with the original user query and the provided dataset description. An algorithm for VisPath is detailed in Algorithm 1 as shown below:
@@
2:// Step 1: Multi-Path Query Expansion
3:$\{R_{\_}1,R_{\_}2,\dots,R_{\_}K\}\leftarrow f_{\_}{\text{mpa}}(Q,D)$
4:// Step 2: Code Generation from Reasoning Paths
5:for $i=1$ to $K$ do
6:     $C_{\_}i\leftarrow f_{\_}{\text{code}}(D,R_{\_}i)$
7:     $V_{\_}i\leftarrow f_{\_}{\text{exec}}(C_{\_}i)$
8:     $\epsilon_{\_}i\leftarrow\begin{cases}1,&\text{if }C_{\_}i\text{ executes successfully}\\ 0,&\text{otherwise}\end{cases}$
9:     $Z_{\_}i\leftarrow\begin{cases}\text{plot image}(V_{\_}i),&\text{if }\epsilon_{\_}i=1\\ \text{error message from }C_{\_}i,&\text{if }\epsilon_{\_}i=0\end{cases}$
10:end for
11:// Step 3: Feedback-Driven Code Optimization
12:for $i=1$ to $K$ do
13:     $F_{\_}i\leftarrow f_{\_}{\text{feedback}}(Q,C_{\_}i,Z_{\_}i)$
14:     $S_{\_}i\leftarrow(C_{\_}i,F_{\_}i)$
15:end for
16:$C^{*}\leftarrow f_{\_}{\text{integrate}}\bigl{(}Q,D,\{S_{\_}1,S_{\_}2,\dots,S_{\_}K\}\bigr{)}$
17:return $C^{*}$
@@ 4 Experiments
4.1 Setup
In this section, we detail our experimental configuration, including (1) experimental datasets, (2) model specifications, and (3) baseline methods for evaluating the performance of the proposed VisPath.
4.1.1 Experimental Datasets
We evaluate our approach on two Text-to-Visualization benchmarks: MatPlotBench Yang et al. (2024) and Qwen-Agent Code Interpreter Benchmark. Specifically, MatPlotBench comprises 100 items with ground truth images; we focus on its simple instruction subset for nuanced queries. In contrast, the Qwen-Agent Code Interpreter Benchmark includes 295 records: 163 related to visualization, and evaluates Python code interpreters on tasks such as mathematical problem solving, data visualization, and file handling based on Code Executability and Code Correctness.
4.1.2 Models Used
Large Language Models (LLMs):
For the code inference stage, we experiment with GPT-4o mini Achiam et al. (2023) and Gemini 2.0 Flash Team et al. (2024) to generate candidate visualization code from the reasoning paths. Both models are configured with a temperature of 0.2 to ensure precise and focused outputs, in line with previous work Yang et al. (2024). To evaluate the generated code quality and guide the subsequent optimization process, we utilize GPT-4o Achiam et al. (2023) and Gemini 2.0 Flash Team et al. (2024) as our visualization feedback model, which provides high-quality reference assessments.
Vision-Language Models (VLMs):
4.1.3 Evaluation Metrics
In our experiments, we utilized evaluation metrics introduced by previous work to ensure consistency and comparability. MatPlotBench Bisong and Bisong (2019) assesses graph generation models using two key metrics: Plot Score, which measures similarity to the Ground Truth (0–100), and Executable Score, which represents the percentage of error-free code executions. Qwen-Agent Code Interpreter benchmark [^2] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/codeinterpreter/README.md evaluates visualization models based on Visualization-Hard and Visualization-Easy. Compared to MatPlotBench, Qwen-Agent Code Interpreter benchmark assesses image alignment via a code correctness metric. Previous studies showed that GPT-based VLM evaluations align well with human assessments Yang et al. (2024), hence VLM was used for evaluation.